OpenStudy (anonymous):

how many photons are produced in a laser pulse of 0.755 J at 513nm?

OpenStudy (aaronq):

first find how much energy 1 photon of that wavelength has, using plancks equation: \(E_{photon}=\dfrac{h*c}{\lambda}\) then divide the total energy (0.755 J) by the energy of 1 photon (that you found).

OpenStudy (anonymous):

Please give me the equation!

OpenStudy (aaronq):

i did, maybe you can't see latex. it's E = h*c/wavelength c =speed of light h=plancks constant

OpenStudy (anonymous):

Yeah...i got it now thanks!

OpenStudy (aaronq):

no probs